# CellFlow: Simulating Cellular Morphology Changes via Flow Matching

### v1

@@ 4.3 Main Results
CellFlow generates highly realistic cell images. CellFlow outperforms existing methods in capturing cellular morphology across all datasets (Table 1a), achieving overall FID scores of 18.7, 33.0, and 9.0 on BBBC021, RxRx1, and JUMP, respectively — improving FID by 21%–45% compared to previous methods. These gains in both FID and KID metrics confirm that CellFlow produces significantly more realistic cell images than prior approaches.
@@
CellFlow accurately captures perturbation-specific morphological changes. As shown in Table 1a, CellFlow achieves conditional FID scores of 56.8 (a 26% improvement), 163.5, and 84.4 (a 16% improvement) on BBBC021, RxRx1, and JUMP, respectively. These scores are computed by measuring the distribution distance for each specific perturbation and averaging across all perturbations. Table 1b further highlights CellFlow’s performance on six representative chemical and three genetic perturbations. For chemical perturbations, CellFlow reduces FID scores by 14–55% compared to prior methods. The smaller improvement (5–12% improvements) on RxRx1 is likely due to the limited number of images per perturbation type.
@@
CellFlow preserves biological fidelity across perturbation conditions. Table 2a presents mode of action (MoA) classification accuracy on the BBBC021 dataset using generated cell images. MoA describes how a drug affects cellular function and can be inferred from morphology. To assess this, we train an image classifier on real perturbed images and test it on generated ones. CellFlow achieves 71.1% MoA accuracy, closely matching real images (72.4%) and significantly surpassing other methods (best: 63.7%), demonstrating its ability to maintain biological fidelity across perturbations. Qualitative comparisons in Figure 3 further highlight CellFlow’s accuracy in capturing key biological effects. For example, demecolcine produces smaller, fragmented nuclei, which other methods fail to reproduce accurately.
@@
CellFlow generalizes to out-of-distribution (OOD) perturbations. On BBBC021, CellFlow demonstrates strong generalization to novel chemical perturbations never seen during training (Table 2b). It achieves 6% and 28% improvements in overall and conditional FID over the best baseline. This OOD generalization is critical for biological research, enabling the exploration of previously untested interventions and the design of new drugs.
@@
Ablations highlight the importance of each component in CellFlow. Table 2c shows that removing conditional information, classifier-free guidance, or noise augmentation significantly degrades performance, leading to higher FID scores. These underscore the critical role of each component in enabling CellFlow’s state-of-the-art performance.
@@
Figure 4: CellFlow enables new capabilities. (a.1) Batch effect calibration. CellFlow initializes with control images, enabling batch-specific predictions. Comparing predictions from different batches highlights actual perturbation effects (smaller cell size) while filtering out spurious batch effects (cell density variations). (a.2) Interpolation trajectory. CellFlow’s learned velocity field supports interpolation between cell states, which might provide insights into the dynamic cell trajectory. (b) Diffusion model comparison. Unlike flow matching, diffusion models that start from noise cannot calibrate batch effects or support interpolation. (c) Reverse trajectory. CellFlow’s reversible velocity field can predict prior cell states from perturbed images, offering potential applications such as restoring damaged cells.
@@ 4.4 New Capabilities
CellFlow addresses batch effects and reveals true perturbation effects. CellFlow’s distribution-to-distribution approach effectively addresses batch effects, a significant challenge in biological experimental data collection. As shown in Figure 4a, when conditioned on two distinct control images with varying cell densities from different batches, CellFlow consistently generates the expected perturbation effect (cell shrinkage due to mevinolin) while recapitulating batch-specific artifacts, revealing the true perturbation effect. Table 2d further quantifies the importance of conditioning on the same batch. By comparing generated images conditioned on control images from the same or different batches against the target perturbation images, we find that same-batch conditioning reduces overall and conditional FID by 21%. This highlights the importance of modeling control images to more accurately capture true perturbation effects—an aspect often overlooked by prior approaches, such as diffusion models that initialize from noise (Figure 4b).
CellFlow has the potential to model cellular morphological change trajectories. Cell trajectories could offer valuable information about perturbation mechanisms, but capturing them with current imaging technologies remains challenging due to their destructive nature. Since CellFlow continuously transforms the source distribution into the target distribution, it can generate smooth interpolation paths between initial and final predicted cell states, producing video-like sequences of cellular transformation based on given source images (Figure 4a). This suggests a possible approach for simulating morphological trajectories during perturbation response, which diffusion methods cannot achieve (Figure 4b). Additionally, the reversible distribution transformation learned through flow matching enables CellFlow to model backward cell state reversion (Figure 4c), which could be useful for studying recovery dynamics and predicting potential treatment outcomes.
@@ 5 Related Works
Cellular morphology prediction. Cellular morphology serves as a powerful phenotypic readout in biological research, offering critical insights into cellular states (Perlman et al., 2004; Loo et al., 2007). Predicting morphological changes in silico enables rapid virtual drug screening and the development of personalized therapeutic strategies, significantly accelerating biomedical discoveries (Carpenter, 2007; Bunne et al., 2024). While initial progress has been made in this direction, existing approaches face three major limitations. Some neglect control cell images, failing to capture true perturbation changes and making predictions vulnerable to batch effects (Yang et al., 2021; Navidi et al., 2024; Cook et al., 2024). Others rely on outdated generative techniques such as normalizing flows and GANs, which suffer from training instability and limited image fidelity (Lamiable et al., 2023; Palma et al., 2023). Additionally, some methods use suboptimal approaches to model distribution transformation, such as a two-step diffusion process (Bourou et al., 2024; Hung et al., 2024). Our work addresses these challenges by reframing morphology prediction as a distribution-to-distribution translation problem and leveraging flow matching, which naturally models cellular state transformations while ensuring high image quality and stable training, paving the way for constructing virtual cells for biomedical research.
@@ 6 Conclusion
In this work, we introduce CellFlow, a method that leverages flow matching to generate cell images under various perturbations while capturing their trajectories, paving the way for the development of a virtual cell framework for biomedical research. In future work, we plan to scale up CellFlow to process terabytes of imaging data encompassing diverse cell types and a wide range of perturbations, enabling the full potential of virtual cell modeling.
@@
CellFlow introduces a novel framework for modeling cellular behavior under genetic and chemical perturbations, which uses flow-matching to generate high-fidelity cellular images and predict phenotypic trajectories. This tool addresses critical challenges in experimental biology by providing scalable and interpretable computational models for studying perturbations at both single-cell and population levels. CellFlow has the potential to accelerate therapeutic discovery and drug repurposing by rapidly screening compounds through in-silico simulations of cell responses to perturbations. Follow-up biology experiments would be directed toward the most promising candidates based on computational experiments. By modeling the space of genetic and chemical perturbations, CellFlow can facilitate the identification of novel therapeutic targets and compounds, streamlining biomedical research. In addition to medical applications, CellFlow can accelerate basic research into cell biology processes by modeling responses to genetic or chemical perturbations. However, we acknowledge that these are early attempts to model complex and dynamic biological systems, and future research with larger and more diverse datasets will improve performance. Furthermore, we are limited by current datasets that focus on a few cancer cell lines, which could introduce bias and may not fully represent normal physiology. We are committed to ensuring robustness in our models and mitigating biases to the extent possible, given the constraints of the dataset and the available training data. In summary, CellFlow bridges machine learning and cellular biology, enabling new frontiers in virtual cell modeling, drug discovery, and systems biology research with broad implications for science and medicine.
@@ Appendix A Theory Proof
Appendix B CellFlow Algorithm
Algorithm 1 CellFlow Algorithm
@@ Appendix B CellFlow Algorithm
Algorithm 1 provides a detailed overview of the CellFlow algorithm, covering both training and inference. During training, the model learns to predict velocity between an initial and target distribution by interpolating between samples, applying noise and condition dropout, and optimizing an L2 loss between predicted and true velocities. In inference, the trained model iteratively updates a sample from the initial distribution toward the target distribution using classifier-free guidance, ultimately generating a new sample that aligns with the target distribution.
@@ Appendix C Experimental Details
Model architecture. CellFlow employs a UNet-based velocity field parameterization with input and output channels matching the dataset. It features four stages for downsampling and upsampling, with each stage halving or doubling the resolution and using a hidden size of 128. This hierarchical UNet design focuses on efficient 2D spatial learning for pixel-level flow matching.
Perturbation encoding. We encode perturbations following IMPA’s approach (Palma et al., 2023). For chemical embeddings, we use 1024-dimensional Morgan Fingerprints generated with RDKit. For gene embeddings, CRISPR and ORF embeddings combine Gene2Vec with HyenaDNA-derived sequence representations, resulting in final dimensions of 328 and 456, respectively.
@@ 3. Visualization of Batch Effects
4. How CellFlow Addresses Batch Effects?
CellFlow mitigates batch effects by using control images as initialization during both training and inference, transporting them to target images within the same batch. This ensures that the model learns only the relative difference between control and perturbed images. By conditioning on control images from different batches, CellFlow effectively captures the true perturbation effect while preserving batch-specific artifacts. Figure 5 demonstrates this, showing that predicted images remain within the same batch cluster when given a control image from that cluster.
Figure 5: Visualization of batch effects in BBBC021 and how CellFlow addresses batch effects.
@@ Appendix G Trajectory
Trajectory examples. Figures 9 and 10 illustrate these bidirectional transformations. The top section of each figure depicts the forward trajectory, where the control image is progressively updated based on the learned velocity field, ultimately generating the perturbed image at $t=1$. The bottom section shows the reverse trajectory, where the process is reversed, progressively reconstructing the original control image. This capability, which is absent in diffusion-based methods, offers a promising approach for simulating morphological trajectories during perturbation responses. Moreover, CellFlow’s reversible distribution transformation enables modeling of backward transitions in cell states, with potential applications in studying recovery dynamics and predicting treatment outcomes.
@@ Appendix I Related Works
IMPA (Palma et al., 2023) GAN Yes Add AdaIn layers to GAN
MorphoDiff (Navidi et al., 2024) Diffusion No -
@@
CellFlow (Ours) Flow Matching Yes Initialization as source distribution

### v3

@@ 4.3 Main Results
CellFlux generates highly realistic cell images. CellFlux outperforms existing methods in capturing cellular morphology across all datasets (Table 1a), achieving overall FID scores of 18.7, 33.0, and 9.0 on BBBC021, RxRx1, and JUMP, respectively — improving FID by 21%–45% compared to previous methods. These gains in both FID and KID metrics confirm that CellFlux produces significantly more realistic cell images than prior approaches.
@@
CellFlux accurately captures perturbation-specific morphological changes. As shown in Table 1a, CellFlux achieves conditional FID scores of 56.8 (a 26% improvement), 163.5, and 84.4 (a 16% improvement) on BBBC021, RxRx1, and JUMP, respectively. These scores are computed by measuring the distribution distance for each specific perturbation and averaging across all perturbations. Table 1b further highlights CellFlux’s performance on six representative chemical and three genetic perturbations. For chemical perturbations, CellFlux reduces FID scores by 14–55% compared to prior methods. The smaller improvement (5–12% improvements) on RxRx1 is likely due to the limited number of images per perturbation type.
@@
CellFlux preserves biological fidelity across perturbation conditions. Table 2a presents mode of action (MoA) classification accuracy and F1 on the BBBC021 dataset using generated cell images. MoA describes how a drug affects cellular function and can be inferred from morphology. To assess this, we train an image classifier on real perturbed images and test it on generated ones. CellFlux achieves 71.1% MoA accuracy, closely matching real images (72.4%) and significantly surpassing other methods (best: 63.7%), demonstrating its ability to maintain biological fidelity across perturbations. Qualitative comparisons in Figure 3 further highlight CellFlux’s accuracy in capturing key biological effects. For example, demecolcine produces smaller, fragmented nuclei, which other methods fail to reproduce accurately.
@@
CellFlux generalizes to out-of-distribution (OOD) perturbations. On BBBC021, CellFlux demonstrates strong generalization to novel chemical perturbations never seen during training (Table 2b). It achieves 6%, 28%, and 170% improvements in overall/conditional FID and MoA accuracy over the best baseline. OOD generalization is critical for biological research, enabling the exploration of previously untested interventions and the design of new drugs.
@@
Ablations highlight the importance of each component in CellFlux. Table 2d shows that removing conditional information, classifier-free guidance, or noise augmentation significantly degrades performance, leading to higher FID scores. These underscore the critical role of each component in enabling CellFlux’s state-of-the-art performance.
@@
Figure 4: CellFlux enables new capabilities. (a.1) Batch effect calibration. CellFlux initializes with control images, enabling batch-specific predictions. Comparing predictions from different batches highlights actual perturbation effects (smaller cell size) while filtering out spurious batch effects (cell density variations). (a.2) Interpolation trajectory. CellFlux’s learned velocity field supports interpolation between cell states, which might provide insights into the dynamic cell trajectory. (b) Diffusion model comparison. Unlike flow matching, diffusion models that start from noise cannot calibrate batch effects or support interpolation. (c) Reverse trajectory. CellFlux’s reversible velocity field can predict prior cell states from perturbed images, offering potential applications such as restoring damaged cells.
@@ 4.4 New Capabilities
CellFlux addresses batch effects and reveals true perturbation effects. CellFlux’s distribution-to-distribution approach effectively addresses batch effects, a significant challenge in biological experimental data collection. As shown in Figure 4a, when conditioned on two distinct control images with varying cell densities from different batches, CellFlux consistently generates the expected perturbation effect (cell shrinkage due to mevinolin) while recapitulating batch-specific artifacts, revealing the true perturbation effect. Table 2c further quantifies the importance of conditioning on the same batch. By comparing generated images conditioned on control images from the same or different batches against the target perturbation images, we find that same-batch conditioning improves conditional FID and MoA accuracy by 14% and 48%. This highlights the importance of modeling control images to more accurately capture true perturbation effects—an aspect often overlooked by prior approaches, such as diffusion models that initialize from noise (Figure 4b).
CellFlux has the potential to model cellular morphological change trajectories. Cell trajectories could offer valuable information about perturbation mechanisms, but capturing them with current imaging technologies remains challenging due to their destructive nature. Since CellFlux continuously transforms the source distribution into the target distribution, it can generate smooth interpolation paths between initial and final predicted cell states, producing video-like sequences of cellular transformation based on given source images (Figure 4a). This suggests a possible approach for simulating morphological trajectories during perturbation response, which diffusion methods cannot achieve (Figure 4b). Additionally, the reversible distribution transformation learned through flow matching enables CellFlux to model backward cell state reversion (Figure 4c), which could be useful for studying recovery dynamics and predicting potential treatment outcomes.
@@ 5 Related Works
Cellular morphology prediction. Cellular morphology serves as a powerful phenotypic readout in biological research, offering critical insights into cellular states (Perlman et al., 2004; Loo et al., 2007). Predicting morphological changes in silico enables rapid virtual drug screening and the development of personalized therapeutic strategies, significantly accelerating biomedical discoveries (Carpenter, 2007; Bunne et al., 2024). While initial progress has been made in this direction, existing approaches face three major limitations. Some neglect control cell images, failing to capture true perturbation changes and making predictions vulnerable to batch effects (Yang et al., 2021; Navidi et al., 2025; Cook et al., 2024). Others rely on outdated generative techniques such as normalizing flows and GANs, which suffer from training instability and limited image fidelity (Lamiable et al., 2023; Palma et al., 2025). Additionally, some methods use suboptimal approaches to model distribution transformation, such as a two-step diffusion process (Bourou et al., 2024; Hung et al., 2024). Our work addresses these challenges by reframing morphology prediction as a distribution-to-distribution translation problem and leveraging flow matching, which naturally models cellular state transformations while ensuring high image quality and stable training, paving the way for constructing virtual cells for biomedical research.
@@ 6 Conclusion
In this work, we introduce CellFlux, a method that leverages flow matching to generate cell images under various perturbations while capturing their trajectories, paving the way for the development of a virtual cell framework for biomedical research. In future work, we plan to scale up CellFlux to process terabytes of imaging data encompassing diverse cell types and a wide range of perturbations, enabling the full potential of virtual cell modeling.
Acknowledgments
This work is partially supported by the Hoffman-Yee Research Grants. E.L. and S.Y. are Chan Zuckerberg Biohub — San Francisco Investigators.
@@ Acknowledgments
CellFlux introduces a novel machine learning framework for modeling cellular responses to genetic and chemical perturbations by formulating the task as a distribution-to-distribution transformation and solving it using a principled flow matching approach. This leads to significantly improved predictive performance and unlocks new capabilities such as batch effect correction and perturbation interpolation.
By providing scalable and interpretable computational tools for modeling perturbation responses at both the single-cell and population levels, CellFlux addresses critical challenges in experimental biology. It enables rapid in-silico screening of compounds and perturbations, thereby accelerating therapeutic discovery and drug repurposing. In particular, it can guide follow-up experiments toward the most promising candidates, streamlining the drug repurposing pipeline and the search for novel therapeutic targets. In addition to medical applications, CellFlux can accelerate basic research into cell biology processes by modeling responses to genetic or chemical perturbations.
However, we acknowledge that these are early attempts to model complex and dynamic biological systems, and future research with larger and more diverse datasets will improve performance. For instance, we are limited by current datasets that focus on a few cancer cell lines, which could introduce bias and may not fully represent normal physiology. Furthermore, while our method enables interpolation between cell states, the biological validity of these interpolations remains unverified; establishing their plausibility will require future work involving ground-truth data and experimental validation.
Despite these limitations, CellFlux bridges machine learning and cellular biology, enabling new frontiers in virtual cell modeling, drug discovery, and systems biology research with broad implications for science and medicine.
@@ Appendix A Theory Proof
Appendix B CellFlux Algorithm
Algorithm 1 CellFlux Algorithm
@@ Appendix B CellFlux Algorithm
Algorithm 1 provides a detailed overview of the CellFlux algorithm, covering both training and inference. During training, the model learns to predict velocity between an initial and target distribution by interpolating between samples, applying noise and condition dropout, and optimizing an L2 loss between predicted and true velocities. In inference, the trained model iteratively updates a sample from the initial distribution toward the target distribution using classifier-free guidance, ultimately generating a new sample that aligns with the target distribution.
@@ Appendix C Experimental Details
Model architecture. CellFlux employs a UNet-based velocity field parameterization with input and output channels matching the dataset. It features four stages for downsampling and upsampling, with each stage halving or doubling the resolution and using a hidden size of 128. This hierarchical UNet design focuses on efficient 2D spatial learning for pixel-level flow matching.
Perturbation encoding. We encode perturbations following IMPA’s approach (Palma et al., 2025). For chemical embeddings, we use 1024-dimensional Morgan Fingerprints generated with RDKit. For gene embeddings, CRISPR and ORF embeddings combine Gene2Vec with HyenaDNA-derived sequence representations, resulting in final dimensions of 328 and 456, respectively.
@@ 3. Visualization of Batch Effects
4. How CellFlux Addresses Batch Effects?
CellFlux mitigates batch effects by using control images as initialization during both training and inference, transporting them to target images within the same batch. This ensures that the model learns only the relative difference between control and perturbed images. By conditioning on control images from different batches, CellFlux effectively captures the true perturbation effect while preserving batch-specific artifacts. Figure 5 demonstrates this, showing that predicted images remain within the same batch cluster when given a control image from that cluster.
Figure 5: Visualization of batch effects in BBBC021 and how CellFlux addresses batch effects.
@@ Appendix G Trajectory
Trajectory examples. Figures 9 and 10 illustrate these bidirectional transformations. The top section of each figure depicts the forward trajectory, where the control image is progressively updated based on the learned velocity field, ultimately generating the perturbed image at $t=1$. The bottom section shows the reverse trajectory, where the process is reversed, progressively reconstructing the original control image. This capability, which is absent in diffusion-based methods, offers a promising approach for simulating morphological trajectories during perturbation responses. Moreover, CellFlux’s reversible distribution transformation enables modeling of backward transitions in cell states, with potential applications in studying recovery dynamics and predicting treatment outcomes.
@@ Appendix H More Results
Effect of sample size on FID/KID. FID and KID are known to be sensitive to the number of samples. We evaluate performance across varying sample sizes on BBBC021 (1K–5K, limited to 6K test images) and JUMP (10K–20K). As shown in Table 5, CellFlux consistently outperforms all baselines across all sample sizes, achieving 30–45% relative improvement and demonstrating the robustness of its improvement regardless of sample size.
Method 1K FID 2.5K FID 5K FID 10K FID 20K FID 1K KID 2.5K KID 5K KID 10K KID 20K KID
PhenDiff 71.3 64.3 49.5 47.5 46.1 2.55 3.68 3.10 4.95 5.09
IMPA 52.4 41.4 33.7 14.0 12.9 3.20 3.38 2.60 1.04 1.05
CellFlux 34.7 25.2 18.7 8.5 7.5 1.67 1.90 1.62 0.63 0.63
Table 5: FID and KID across different sample sizes.
Comparison with more baselines. Cell morphology prediction is a new task with only six baselines (Table 8). We included the only two published methods using control images (Bourou et al., 2024; Palma et al., 2025); others are unpublished (Hung et al., 2024; Cook et al., 2024), lack code (Yang et al., 2021; Cook et al., 2024), or omit controls (Yang et al., 2021; Navidi et al., 2025; Cook et al., 2024). We further compared MorphoDiff (Navidi et al., 2025), a recent diffusion-based method, without using control images. Under our setup on BBBC021, CellFlux outperforms it in image quality and MoA metrics.
Method FIDo FIDc KIDo KIDc MoA Accuracy MoA Macro-F1 MoA Weighted-F1
MorphoDiff 65.8 114.1 7.99 7.97 38.3 24.5 34.2
CellFlux 18.7 56.8 1.62 1.59 71.2 49.0 70.7
Table 6: CellFlux outperforms MorphoDiff on BBBC021 in both image quality and MoA classification metrics.
Cross-dataset transfer. To assess whether the learned model can generalize to highly out-of-distribution settings, we conduct a cross-dataset transfer experiment by applying a CellFlux model trained on BBBC021 to RxRx1 and JUMP images. Surprisingly, we observe that CellFlux successfully transfers and applies perturbation effects despite substantial domain shifts (Figure 11), highlighting its potential as a unified foundation model across diverse perturbation datasets. Note that there are no target images for RxRx1 and JUMP, as these datasets do not contain the corresponding perturbation conditions.
Figure 11: Cross-dataset transfer of CellFlux. Although CellFlux is trained solely on BBBC021, it demonstrates zero-shot generalization to two unseen datasets—RxRx1 and JUMP. Notably, it can predict morphological changes induced by perturbations (AZ138 and Demecolcine) that are absent from both datasets. AZ138, an Eg5 inhibitor, leads to cell shrinkage and death, while Demecolcine disrupts microtubules, resulting in smaller, fragmented nuclei.
CellProfiler metrics. To further evaluate whether CellFlux can capture perturbation-specific morphological changes, we extracted CellProfiler features related to nuclear size under three perturbations known to enlarge nuclei (taxol, vincristine, and demecolcine) using the BBBC021 dataset. As shown in Table 7 (mean and 95% confidence interval reported), CellFlux most closely matches the real perturbed morphology in terms of nuclear size.
Method taxol vincristine demecolcine
Control 1612.0 ± 39.5 1612.0 ± 39.5 1612.0 ± 39.5
Target 2296.7 ± 190.1 2365.5 ± 125.5 2311.0 ± 136.1
PhenDiff 1755.9 ± 138.2 1947.8 ± 70.8 2118.8 ± 102.7
IMPA 2088.3 ± 190.8 2116.9 ± 107.1 2386.5 ± 123.9
CellFlux 2141.0 ± 166.6 2276.4 ± 115.6 2323.8 ± 121.9
Table 7: Comparison of CellProfiler nuclear size features under three compounds known to enlarge nuclei.
@@ Appendix I Related Works
MorphoDiff (Navidi et al., 2025) Diffusion No -
IMPA (Palma et al., 2025) GAN Yes Add AdaIn layers to GAN
CellFlux (Ours) Flow Matching Yes Initialization as source distribution
@@
Appendix J Biological Validation of Interpolation Trajectories
While CellFlux enables interpolation of perturbation trajectories, the biological relevance of the interpolated states remains unverified. Validating these trajectories is a key next step. Although existing datasets lack ground-truth labels for intermediate states, future work could explore the following directions:
  • Dose interpolation: Some datasets include images under multiple dosage levels. We can test whether interpolation from control to high dose yields intermediate states that resemble realistic medium-dose morphologies.
  • Timepoint interpolation: For datasets with multiple timepoints (e.g., day 0, day 5, day 10), we can evaluate whether interpolation from control to day 10 recovers morphologies consistent with intermediate timepoints such as day 5.
  • Drug perturbation interpolation: Current datasets rarely include fine-grained trajectories post-treatment. Validating such interpolations may require collecting new wet-lab data, such as live-cell imaging over time.
In addition to validation, future work could improve the biological plausibility of interpolation trajectories and avoid degenerate “shortest path in pixel space” artifacts by:
  • Interpolating in latent space: Performing interpolation in a learned latent space (e.g., via an autoencoder) rather than pixel space may encourage trajectories to follow a more structured biological manifold.
  • Adding supervision from intermediate states: When datasets contain known intermediate states (e.g., medium dose), models can be trained to explicitly pass through these points during interpolation.
  • Adding constraints: Incorporating additional constraints—such as a GAN loss—can guide interpolated images to resemble real cell images, improving biological realism.